# Can We Treat Noisy Labels as Accurate?

### v1

@@ Appendix D Runtime Analysis
The efficiency of EchoMod is significantly influenced by several factors, including the choice of the controllable generative model, the GPU’s floating-point operations per second (FLOPS), the resolution of generated images, batch size, GPU memory capacity, floating-point precision, and the number of diffusion steps if a diffusion model is used. In this study, we employ an NVIDIA V100-SXM2 with 32GB of VRAM, using ControlNet as the benchmark generative model, and apply mixed precision to assess the impacts of image size and batch size on runtime. Runtime is measured in GPU hours, which quantifies the time consumed to perform computational tasks on a single GPU. As the number of GPUs increases, we observe a super-linear decrease in runtime. Our experiments are conducted on the CIFAR-10 dataset, and we also estimate the runtime for processing the Clothing1M dataset on the same GPU configuration. Table 6 demonstrates that increasing the batch size and reducing the image resolution both significantly impact runtime. We did not conduct tests with image resolution at $768\times 768$ and a batch size of 16 due to similar VRAM constraints. Additionally, in Table 7, we compare the effects of different computing performance on runtime. We conducted tests on two different GPUs with an image resolution of $512\times 512$ and a batch size of 8. The NVIDIA V100-SXM2-32GB offers a half-precision compute capability of 125 Tensor TFLOPS and a single-precision capability of 15.7 TFLOPS. In contrast, the more powerful NVIDIA RTX 4090-24GB GPU provides 165.2 Tensor TFLOPS in half-precision and 82.58 TFLOPS in single-precision.
Appendix E Challenges of Instance Modification
Instance modification presents several significant challenges that need to be addressed to ensure the effectiveness and reliability of the modified instances. Instance modification primarily relies on Controllable Generative Models (CGM), which can generate instances that meet specific requirements under certain conditions. However, Characteristic Shift is a major challenge, as illustrated in Figure 7(a). Characteristic Shift occurs during instance modification when the fundamental attributes of the instances are altered. These changes can lead to inconsistencies and distortions in the original data, the effect majorly due to the gap between the intrinsic features of the original instances and the noise labels. For example, if the CGMs significantly alter key features such as shape or texture during modification, the models may fail to learn consistent representations. Therefore, addressing the Characteristic Shift between original instances and noise labels is crucial. Another major challenge is Distribution Shift, as shown in Figures 7(b) and 7(c). This challenge arises when the statistical distribution of the modified instances deviates from the original data distribution. Such shifts can lead to models performing well on modified data but poorly on real-world data. The T-SNE visualization of CIFAR-10 instance representations shows the changes in instance distribution before and after using EchoAlign, showcasing the effectiveness of our method.
(a) Characteristic Shift.
Appendix F More Generation Examples
In this section, we present the results of several publicly available controllable (CGM) and non-controllable (NGM) generative models that can easily generate demos. We used “a photo of noisy label” as the prompt and the original instance as the control condition. As shown in Table 8, for controllable generative models, we present the results of ControlNet and UniControl [Zhao et al., 2023]. For non-controllable generative models, we showcase the performance of GPT-4 [Achiam et al., 2023] and Gemini [Team et al., 2023].
CGM NGM
Noisy Label Original Instance ControlNet UniControl GPT-4 Gemini
Appendix G Framework
Figure 8 illustrates the framework of EchoAlign, which comprises two main modules: EchoMod and EchoSelect. The EchoMod module uses Controllable Generative Models (CGM) to modify noisy instances, generating instances consistent with their labels. Specifically, the CGM receives noisy instances and generates new instances based on the images’ labels and intrinsic features. Through the EchoMod module, we obtain a set of modified instances. Subsequently, these modified instances enter the EchoSelect module. In this module, the system uses cosine similarity evaluation to select the instances from the noisy data that are most similar to the modified instances. The similarity evaluation assumes that the features of clean instances are highly consistent with those of the modified instances, thereby selecting sufficiently clean samples to enhance the model’s generalization capability. Finally, the selected clean instances and the modified instances together form the training instance set used for subsequent model training. In this way, the EchoAlign framework effectively addresses the issues of characteristic shift and distribution shift in instance modification through its robust instance modification and selection processes, ensuring the model’s generalization ability and robustness.
Figure 8: The framework of EchoAlign.

### v2

@@ 5.4 Runtime Analysis
The efficiency of EchoMod is significantly influenced by several factors, including the choice of the controllable generative model, the GPU’s floating-point operations per second (FLOPS), the resolution of generated images, batch size, GPU memory capacity, floating-point precision, and the number of diffusion steps if a diffusion model is used. In this study, we employ an NVIDIA V100-SXM2 with 32GB of VRAM, using ControlNet as the benchmark generative model, and apply mixed precision to assess the impacts of image size and batch size on runtime. Runtime is measured in GPU hours, representing the computational time required to perform tasks on a single GPU. As the number of GPUs increases, we observe a super-linear reduction in runtime. Our experiments are conducted on the CIFAR-10 dataset, and we also estimate the runtime for processing the Clothing1M dataset on the same GPU configuration. Table 7 demonstrates that increasing the batch size and reducing the image resolution both significantly impact runtime. We did not conduct tests with image resolution at $768\times 768$ and a batch size of 16 due to GPU memory constraints. Additionally, in Table 8, we compare the effects of different computing performance on runtime. We conducted tests on two different GPUs with an image resolution of $512\times 512$ and a batch size of 8. The NVIDIA V100-SXM2-32GB offers a half-precision compute capability of 125 Tensor TFLOPS and a single-precision capability of 15.7 TFLOPS. In contrast, the more powerful NVIDIA RTX 4090-24GB GPU provides 165.2 Tensor TFLOPS in half-precision and 82.58 TFLOPS in single-precision. Although using ControlNet as a benchmark for evaluating the Clothing1M dataset is not optimal in terms of inference efficiency, the high flexibility of our framework allows for substantial improvements by incorporating various optimization techniques and more advanced models. For instance, employing optimization strategies from Ultra Fast ControlNet (Paul et al.,, 2023), such as efficient schedulers and smart CPU memory offloading, can achieve approximately a 70.59% improvement in inference speed. Moreover, adopting more efficient model architectures, such as ControlNet-XS (Zavadski et al.,, 2024), can significantly enhance inference speed, with reported improvements of up to 46.48%. Based on a preliminary estimation and assuming the use of an RTX 4090-24GB GPU, the inference time on Clothing1M could theoretically be reduced from 338.2 GPU hours to 53.2 GPU hours, even without considering additional memory optimizations that might allow for larger batch sizes.
6 Conclusion
This work provided a novel perspective on treating noisy labels as accurate through instance modification. Theoretical analysis supports that this alignment process allows models to learn meaningful patterns despite the presence of labeling errors. To address the challenges of instance modification, we proposed the EchoAlign framework, which integrates a controllable generative model with strategic sample selection to create a robust training dataset. Extensive experiments on diverse datasets demonstrate the superiority of EchoAlign over existing methods, particularly in scenarios with high levels of label noise. However, EchoAlign’s success partially depends on the capabilities of the controllable generative model, and the cost of fine-tuning these models to adapt to out-of-distribution data could be a barrier in resource-limited settings. Future directions include investigating supervised or self-supervised extensions and broader applications such as medical imaging or real-time systems.
Acknowledgements
This work is supported by the National Natural Science Foundation of China [U23A20389, 62176139], the Major Basic Research Project of the Natural Science Foundation of Shandong Province [ZR2021ZD15]. The author would like to express sincere gratitude to Prof. Tongliang Liu for his invaluable guidance and support throughout this research.
1. Alignment:
[…]
Definitions and Assumptions:
Let $X\in\mathbb{R}^{d}$ be the original instances with distribution $P_{X}$ and $\tilde{Y}\in\mathcal{Y}$ be the noisy labels, where $\mathcal{Y}$ is the label space. The modified instances are defined as $X^{\prime}=T(X,\tilde{Y};\theta)\in\mathbb{R}^{d}$, where $T$ is a transformation function parameterized by $\theta$, designed to improve alignment between $X^{\prime}$ and $\tilde{Y}$.
We make the following assumptions:
  • 1.
    A1. Transformation Improves Alignment: The transformation $T$ reduces the conditional entropy of $\tilde{Y}$ given the features:
    $H(\tilde{Y}\mid X^{\prime})\leq H(\tilde{Y}\mid X).$
    This means that the uncertainty in $\tilde{Y}$ given $X^{\prime}$ is less than or equal to the uncertainty given $X$.
Goal:
Our aim is to prove that the mutual information between $X^{\prime}$ and $\tilde{Y}$ is greater than or equal to that between $X$ and $\tilde{Y}$:
[…]
where the mutual information is defined as:
$I(X;\tilde{Y})=H(\tilde{Y})-H(\tilde{Y}\mid X).$
Proof.
We begin by expressing the mutual information between $X$ (or $X^{\prime}$) and $\tilde{Y}$:
$I(X;\tilde{Y})=H(\tilde{Y})-H(\tilde{Y}\mid X),\quad I(X^{\prime};\tilde{Y})=H(\tilde{Y})-H(\tilde{Y}\mid X^{\prime}).$
The difference in mutual information is then:
[…]
$=\left[H(\tilde{Y})-H(\tilde{Y}\mid X^{\prime})\right]-\left[H(\tilde{Y})-H(\tilde{Y}\mid X)\right]$
$=H(\tilde{Y}\mid X)-H(\tilde{Y}\mid X^{\prime}).$
According to Assumption A1, the transformation $T$ reduces the conditional entropy of $\tilde{Y}$ given the features, so:
$H(\tilde{Y}\mid X^{\prime})\leq H(\tilde{Y}\mid X).$
Therefore, the difference $\Delta I$ is non-negative:
$\Delta I=H(\tilde{Y}\mid X)-H(\tilde{Y}\mid X^{\prime})\geq 0.$
This implies that:
[…]
Thus, the mutual information between the modified instances $X^{\prime}$ and the noisy labels $\tilde{Y}$ is greater than or equal to that between the original instances $X$ and $\tilde{Y}$, indicating improved alignment between $X^{\prime}$ and $\tilde{Y}$.
2. Error Reduction:
Definitions and Assumptions:
Let $X\in\mathbb{R}^{d}$ be the original instances with distribution $P_{X}$, $\tilde{Y}\in\mathbb{R}$ be the noisy labels, and $Y\in\mathbb{R}$ be the true labels. The modified instances are defined as $X^{\prime}=T(X,\tilde{Y};\theta)\in\mathbb{R}^{d}$, where $T$ is a transformation designed to improve alignment between $X^{\prime}$ and $\tilde{Y}$ while preserving essential predictive information about $Y$.
We consider two models:
  • 1.
[…]
  • 2.
[…]
The loss function $L:\mathbb{R}\times\mathbb{R}\rightarrow\mathbb{R}_{\geq 0}$ is assumed to be convex and differentiable with respect to its second argument (e.g., squared loss $L(y,\hat{y})=(y-\hat{y})^{2}$).
We make the following assumptions:
  • 1.
    A1. Transformation Improves Alignment: The transformation $T$ reduces the variance of the noisy labels conditioned on the features:
    $\operatorname{Var}(\tilde{Y}\mid X^{\prime})\leq\operatorname{Var}(\tilde{Y}\mid X).$
  • 2.
    A2. Transformation Preserves Predictive Information: The transformation $T$ preserves the essential information needed to predict $Y$:
    $\operatorname{Var}(Y\mid X^{\prime})\approx\operatorname{Var}(Y\mid X).$
Goal:
Our aim is to prove that the expected risk of $f_{X^{\prime}}$ evaluated on the original instances $X$ is less than or equal to that of $f_{X}$:
$R(f_{X^{\prime}})\leq R(f_{X}),$
where the expected risks are defined as:
$R(f_{X})=\mathbb{E}_{X,Y}\left[L(Y,f_{X}(X))\right],\quad R(f_{X^{\prime}})=\mathbb{E}_{X,Y}\left[L(Y,f_{X^{\prime}}(X))\right].$
Proof.
We decompose the expected risk into the Bayes risk and the excess risk. Let $f^{*}(X)=\mathbb{E}[Y\mid X]$ be the Bayes optimal predictor, which minimizes the expected loss:
$R^{*}=\mathbb{E}_{X,Y}\left[L(Y,f^{*}(X))\right].$
The excess risks for $f_{X}$ and $f_{X^{\prime}}$ are then:
$\mathcal{E}(f_{X})=R(f_{X})-R^{*},\quad\mathcal{E}(f_{X^{\prime}})=R(f_{X^{\prime}})-R^{*}.$
Our goal is to show that $\mathcal{E}(f_{X^{\prime}})\leq\mathcal{E}(f_{X})$.
For each model, the excess risk can be expressed as:
$\mathcal{E}(f)=\mathbb{E}_{X}\left[\mathbb{E}_{Y\mid X}\left[L(Y,f(X))-L(Y,f^{*}(X))\right]\right].$
Assuming $L$ is twice differentiable, we perform a Taylor expansion of $L(Y,f(X))$ around $f^{*}(X)$:
$L(Y,f(X))\approx$ $L(Y,f^{*}(X))+L^{(1)}(Y,f^{*}(X))\delta(X)$
$+\frac{1}{2}L^{(2)}(Y,f^{*}(X))\delta(X)^{2}.$
where $\delta(X)=f(X)-f^{*}(X)$, and $L^{(1)}$, $L^{(2)}$ are the first and second derivatives with respect to the second argument. Since $f^{*}(X)$ minimizes $\mathbb{E}_{Y\mid X}[L(Y,f(X))]$, the expected first derivative term is zero:
$\mathbb{E}_{Y\mid X}[L^{(1)}(Y,f^{*}(X))]=0.$
Thus, the excess risk simplifies to:
$\mathcal{E}(f)\approx\mathbb{E}_{X}\left[\frac{1}{2}\mathbb{E}_{Y\mid X}[L^{(2)}(Y,f^{*}(X))]\delta(X)^{2}\right].$
We focus on comparing $\delta_{X}(X)^{2}$ and $\delta_{X^{\prime}}(X)^{2}$, where $\delta_{X}(X)=f_{X}(X)-f^{*}(X)$ and $\delta_{X^{\prime}}(X)=f_{X^{\prime}}(X)-f^{*}(X)$.
[…]
$\operatorname{Var}(\delta_{X^{\prime}}(X))\leq\operatorname{Var}(\delta_{X}(X)).$
Assuming the biases $\mathbb{E}_{X}[\delta_{X}(X)]$ and $\mathbb{E}_{X}[\delta_{X^{\prime}}(X)]$ are negligible or similar due to Assumption A2, we have:
$\mathbb{E}_{X}[\delta_{X^{\prime}}(X)^{2}]$ $=\operatorname{Var}(\delta_{X^{\prime}}(X))+\left(\mathbb{E}_{X}[\delta_{X^{\prime}}(X)]\right)^{2}$
$\leq\operatorname{Var}(\delta_{X}(X))+\left(\mathbb{E}_{X}[\delta_{X}(X)]\right)^{2}=\mathbb{E}_{X}[\delta_{X}(X)^{2}].$
Since $L^{(2)}$ is positive due to the convexity of $L$, it follows that:
$\mathcal{E}(f_{X^{\prime}})\leq\mathcal{E}(f_{X}).$
Adding back the Bayes risk $R^{*}$, we conclude:
$R(f_{X^{\prime}})=R^{*}+\mathcal{E}(f_{X^{\prime}})\leq R^{*}+\mathcal{E}(f_{X})=R(f_{X}).$
Therefore, the expected prediction error is minimized when using the model trained on the modified instances $X^{\prime}$, even when evaluated on the original data $X$. ∎
3. Estimation Stability:
  • 1.
  • 2.
Goal: To demonstrate that the variance of the estimator $f_{X^{\prime}}$ is lower than that of $f_{X}$.
Proof.
  • 1.
  • 2.
    Variance of Estimators: The covariance of the estimated coefficients under OLS can be expressed as:
    $\text{Cov}(\hat{\beta}_{X})$ $=\sigma^{2}(X^{T}X)^{-1}$
    $\text{Cov}(\hat{\beta}_{X^{\prime}})$ $=\sigma^{2}(X^{\prime T}X^{\prime})^{-1}$
    where $\sigma^{2}$ is the variance of the error terms $\epsilon_{X}$ and $\epsilon_{X^{\prime}}$, assumed equal for simplicity. The variance of the predicted values at any input $x$ and its modified version $x^{\prime}$ is:
    $\text{Var}(f_{X}(x))$ $=x^{T}\text{Cov}(\hat{\beta}_{X})x=\sigma^{2}x^{T}(X^{T}X)^{-1}x$
    $\text{Var}(f_{X^{\prime}}(x^{\prime}))$ $=x^{\prime T}\text{Cov}(\hat{\beta}_{X^{\prime}})x^{\prime}=\sigma^{2}x^{\prime T}(X^{\prime T}X^{\prime})^{-1}x^{\prime}$
  • 3.
    Comparative Analysis of Variance: Since $X^{\prime}$ is designed to be more informative and aligned with $\tilde{Y}$, it is reasonable to assume that $X^{\prime}$ exhibits higher effective variability in the dimensions that are most relevant for predicting $Y$. This increased effective variability implies that the matrix $X^{\prime T}X^{\prime}$ is larger than $X^{T}X$ in the Loewner partial ordering, meaning:
    $X^{\prime T}X^{\prime}\succeq X^{T}X$
    This leads to:
    $(X^{\prime T}X^{\prime})^{-1}\preceq(X^{T}X)^{-1}$
    Because the inverse of a larger positive definite matrix is smaller in the Loewner ordering. Consequently, the covariance matrices satisfy:
    $\text{Cov}(\hat{\beta}_{X^{\prime}})\preceq\text{Cov}(\hat{\beta}_{X})$
    Assuming that the transformation from $x$ to $x^{\prime}$ does not significantly increase the norm of the input vectors (i.e., $\|x^{\prime}\|\approx\|x\|$), we can compare the variances of the predictions:
    $\text{Var}(f_{X^{\prime}}(x^{\prime}))$ $=x^{\prime T}\text{Cov}(\hat{\beta}_{X^{\prime}})x^{\prime}\leq x^{\prime T}\text{Cov}(\hat{\beta}_{X})x^{\prime}$
    $\approx x^{T}\text{Cov}(\hat{\beta}_{X})x=\text{Var}(f_{X}(x))$
    Thus:
    $\text{Var}(f_{X^{\prime}}(x^{\prime}))\leq\text{Var}(f_{X}(x))$
  • 4.
    Estimation Stability: Thus, the variance of the predictions using $X^{\prime}$ is less than or equal to that using $X$. The reduction in variance implies that $f_{X^{\prime}}$ offers more stable and reliable predictions compared to $f_{X}$. This stability is crucial when the model is applied in practice, particularly in the presence of noisy data conditions. This result holds under the assumptions that:
    1. (a)
      $X^{\prime T}X^{\prime}\succeq X^{T}X$ (i.e., $X^{\prime T}X^{\prime}-X^{T}X$ is positive semidefinite).
    2. (b)
      The transformation from $x$ to $x^{\prime}$ does not significantly increase the input vector norms.
4. Generalization:
Setup: Let
  • 1.
    $X$ denote the original feature space and $X^{\prime}=T(X,\tilde{Y};\theta)$ denote the modified feature space, where $T$ is a transformation (assumed to be Lipschitz continuous with Lipschitz constant $L\leq 1$), and $\theta$ is fixed, that optimizes some aspect of the data to better align with noisy labels $\tilde{Y}$. In our EchoAlign framework, the transformation $T$ is implemented using controllable generative models. These models can be designed to be Lipschitz continuous by incorporating techniques like spectral normalization or gradient penalties. Ensuring that $L\leq 1$ is reasonable because we aim for $T$ to be non-expansive, preventing the amplification of noise and promoting stability in the transformation.
  • 2.
Generalization Bounds: Using these definitions, the generalization bounds for a Lipschitz continuous loss function $l$ can be expressed for both feature sets. Assuming the same hypothesis class $\mathcal{F}$, the bounds are:
$\mathbb{E}[l(f(X),Y)]$ $\leq\frac{1}{n}\sum_{i=1}^{n}l(f(X_{i}),Y_{i})+2\mathfrak{R}_{n}(\mathcal{F}_{X})+c$
$\mathbb{E}[l(f(X^{\prime}),Y)]$ $\leq\frac{1}{n}\sum_{i=1}^{n}l(f(X^{\prime}_{i}),Y_{i})+2\mathfrak{R}_{n}(\mathcal{F}_{X^{\prime}})+c$
where $c$ is a constant that depends on the complexity of the loss function.
  • 1.
    Increase the signal-to-noise ratio in $X^{\prime}$ compared to $X$.
  • 2.
Proof.
To show that $\mathfrak{R}_{n}(\mathcal{F}_{X^{\prime}})\leq L\cdot\mathfrak{R}_{n}(\mathcal{F}_{X})$ and hence that $\mathfrak{R}_{n}(\mathcal{F}_{X^{\prime}})\leq\mathfrak{R}_{n}(\mathcal{F}_{X})$ when $L\leq 1$, we analyze how the transformation $T$ affects the ability of the function class $\mathcal{F}$ to fit random noise.
Since $T$ is Lipschitz continuous with Lipschitz constant $L\leq 1$, we can apply the contraction principle (Ledoux-Talagrand contraction inequality) to relate the Rademacher complexities:
$\mathfrak{R}_{n}(\mathcal{F}_{X^{\prime}})=\mathfrak{R}_{n}(\mathcal{F}\circ T)\leq L\cdot\mathfrak{R}_{n}(\mathcal{F}_{X})$
Since $X^{\prime}=T(X,\tilde{Y};\theta)$ and $T$ is Lipschitz continuous in $X$ (with $\tilde{Y}$ and $\theta$ fixed during transformation), the inequality applies directly. When $L\leq 1$, this inequality implies that the Rademacher complexity on the modified data $X^{\prime}$ is less than or equal to that on the original data $X$:
$\mathfrak{R}_{n}(\mathcal{F}_{X^{\prime}})\leq\mathfrak{R}_{n}(\mathcal{F}_{X})$
Since the generalization error bound depends on the Rademacher complexity, a lower Rademacher complexity implies a tighter generalization bound. Specifically:
$\mathbb{E}[l(f(X^{\prime}),Y)]\leq\frac{1}{n}\sum_{i=1}^{n}l(f(X^{\prime}_{i}),Y_{i})+2\mathfrak{R}_{n}(\mathcal{F}_{X^{\prime}})+c$
With $\mathfrak{R}_{n}(\mathcal{F}_{X^{\prime}})\leq\mathfrak{R}_{n}(\mathcal{F}_{X})$, the bound on the expected loss for $X^{\prime}$ is tighter than that for $X$. Therefore, the model trained on $X^{\prime}$ is expected to generalize better than the model trained on $X$.
This inequality derived from comparing the Rademacher complexities and the corresponding generalization bounds provides a theoretical basis for asserting that instance modification enhances the model’s ability to generalize. By ensuring that the transformation $T$ is Lipschitz continuous with $L\leq 1$, we have formally shown that the Rademacher complexity decreases or remains the same, leading to improved generalization performance. This proof underscores the importance of feature alignment and relevance in improving machine learning model performance in noisy settings. ∎
Proof of the Lipschitz Continuity of the Transformation $T$ with $L\leq 1$:
In our framework, the transformation $T$ modifies an original instance $X$ to a modified instance $X^{\prime}=T(X,\tilde{Y};\theta)$, aiming to align $X^{\prime}$ more closely with its noisy label $\tilde{Y}$ while preserving essential characteristics of $X$. To demonstrate that $T$ is Lipschitz continuous with Lipschitz constant $L\leq 1$, we proceed as follows.
We define the transformation $T$ as a convex combination of the original instance $X$ and an adjustment function $\phi(X,\tilde{Y};\theta)$ that incorporates the influence of the noisy label:
$T(X,\tilde{Y};\theta)=(1-\alpha)X+\alpha\phi(X,\tilde{Y};\theta),$
where $\alpha\in[0,1]$ is a parameter controlling the degree of modification, and $\phi(X,\tilde{Y};\theta)$ is designed to adjust $X$ based on $\tilde{Y}$.
To ensure that $T$ is Lipschitz continuous with $L\leq 1$, we require that $\phi$ itself is Lipschitz continuous with Lipschitz constant $L_{\phi}\leq 1$. Under this condition, for any two instances $X_{1},X_{2}\in\mathcal{X}$, we have:
$\|T(X_{1},\tilde{Y};\theta)-T(X_{2},\tilde{Y};\theta)\|$
$=\left\|(1-\alpha)(X_{1}-X_{2})+\alpha\left(\phi(X_{1},\tilde{Y};\theta)-\phi(X_{2},\tilde{Y};\theta)\right)\right\|$
$\leq(1-\alpha)\|X_{1}-X_{2}\|+\alpha\|\phi(X_{1},\tilde{Y};\theta)-\phi(X_{2},\tilde{Y};\theta)\|$
$\leq(1-\alpha)\|X_{1}-X_{2}\|+\alpha L_{\phi}\|X_{1}-X_{2}\|$
$=\left((1-\alpha)+\alpha L_{\phi}\right)\|X_{1}-X_{2}\|.$
Since $L_{\phi}\leq 1$ and $\alpha\in[0,1]$, we have:
$(1-\alpha)+\alpha L_{\phi}\leq(1-\alpha)+\alpha=1,$
which means that the Lipschitz constant $L$ of $T$ satisfies $L\leq 1$.
To ensure that $\phi$ has Lipschitz constant $L_{\phi}\leq 1$, we can design $\phi$ using various techniques:
  • 1.
    Spectral Normalization: Spectral normalization constrains the spectral norm (largest singular value) of each linear layer in the neural network implementing $\phi$ to be at most 1 (Miyato et al.,, 2018). By normalizing the weight matrices $W$ of the layers such that:
    $\|W\|_{2}=\sigma_{\text{max}}(W)=1,$
    we ensure that the Lipschitz constant of each layer does not exceed 1. Since the Lipschitz constant of a composition of functions is bounded by the product of the individual Lipschitz constants, and each layer has $L_{i}\leq 1$, the overall Lipschitz constant of $\phi$ satisfies $L_{\phi}\leq 1$.
  • 2.
    Gradient Penalties: Incorporating gradient penalties into the training of $\phi$ encourages the network to have controlled Lipschitz continuity (Gulrajani et al.,, 2017). We add a regularization term to the loss function:
    $\mathcal{L}_{\text{GP}}=\lambda,\mathbb{E}_{X,\tilde{Y}}\left[\left(\left\|\nabla_{X}\phi(X,\tilde{Y};\theta)\right\|_{2}-1\right)^{2}\right],$
    where $\lambda>0$ is a penalty coefficient. Minimizing $\mathcal{L}_{\text{GP}}$ enforces the gradient norms of $\phi$ to be close to 1, ensuring $L_{\phi}\leq 1$.
  • 3.
    Contractive Autoencoders: Designing $\phi$ as a contractive autoencoder (Rifai et al.,, 2011) involves adding a contraction penalty to the loss function:
    $\mathcal{L}_{\text{CAE}}=\mathbb{E}_{X}\left[\|X-\phi(X,\tilde{Y};\theta)\|_{2}^{2}+\lambda\ \left\|\frac{\partial\phi(X,\tilde{Y};\theta)}{\partial X}\right\|_{F}^{2}\right],$
    where $\|\cdot\|_{F}$ denotes the Frobenius norm, and $\lambda>0$ controls the penalty strength. This penalizes large derivatives, encouraging $\phi$ to be contractive and thus Lipschitz continuous with $L_{\phi}\leq 1$.